# Word Embedding Dimension Reduction via Weakly-Supervised Feature Selection

### v1

@@ Abstract
As a fundamental task in natural language processing, word embedding converts each word into a representation in a vector space. A challenge with word embedding is that as the vocabulary grows, the vector space’s dimension increases and it can lead to a vast model size. Storing and processing word vectors are resource-demanding, especially for mobile edge-devices applications. This paper explores word embedding dimension reduction. To balance computational costs and performance, we propose an efficient and effective weakly-supervised feature selection method, named WordFS. It has two variants, each utilizing novel criteria for feature selection. Experiments conducted on various tasks (e.g., word and sentence similarity and binary and multi-class classification) indicate that the proposed WordFS model outperforms other dimension reduction methods at lower computational costs.
@@ 1 Introduction
In recent years, large language models (LLMs) have made a breakthrough in natural language processing (NLP). These models have revolutionized the understanding and generation of human language tasks such as question-answering, machine translation, and text summarization [28, 29, 16], etc. As a fundamental component in language models, word embedding represents words as vectors in a continuous, high-dimensional space [27]. Word vectors capture semantic and syntactic meanings of words, providing word relationships for various downstream tasks. Static word embedding methods, such as Glove [17], Word2Vec [13], and Fasttext [2], assign a fixed vector to each word. They convert input words to corresponding vectors for further processing and model training. Contextual word embedding methods, such as ELMo [18] and BERT [4], leverage deep-learning models to generate vectors for words based on the context. A word can have a different contextual word embedding in a different sentence, allowing the model to capture subtle differences in the meaning of the same word in various contexts.
@@
Traditional unsupervised PCA-based methods are known for their efficiency and interpretability. In contrast, supervised deep-learning-based methods are inefficient and lack interpretability. For instance, autoencoders take much longer training and inference time in sentence embedding dimension reduction than traditional methods [31]. Semi-supervised feature selection methods can also be used for dimension reduction. Since word vectors can be viewed as extracted features for each word, feature selection methods have the potential to provide a straightforward yet effective way to reduce the word dimension. It could help balance unsupervised PCA-based methods and resource-intensive deep-learning-based methods.
This paper investigates using a small subset of labeled word similarity pairs to develop a weakly supervised dimension reduction method called WordFS (i.e., word dimension reduction with Feature Selection). We demonstrate that one can achieve dimension reduction effectively by supervising a limited number of word similarity pairs. Note that word similarity is typically used in evaluation tasks but rare in word embedding dimension reduction - a key novelty of this work. The proposed WordFS method consists of three stages: 1) post-processing, 2) feature extraction, and 3) weakly-supervised feature selection. We apply WordFS to other downstream tasks to show its generalizability. Experimental results show that WordFS outperforms existing methods in word similarity and various tasks while achieving much lower computational costs.
@@
  • We show the effectiveness and efficiency of our approach on various downstream tasks, including sentence similarity and classification tasks. Our method generally outperforms the existing techniques while being more straightforward.
@@ 2.1 Matrix Decomposition Techniques
Post-processing methods are helpful in enhancing word embeddings in dimension reduction. Effective construction of lower dimensional word embeddings can be achieved by combining post-processing algorithms (PPAs) based on reducing the anisotropy [15] with the PCA method [19]. The dimension-reduced vectors can perform similarly or even better than the original pre-trained word embeddings, outperforming most unsupervised methods. [9] extends this method to contextual embeddings by adding in additional geodesic distance information via the Isomap algorithm [24]. PCA also performs well as an unsupervised dimension reduction method for pre-trained sentence embeddings [31].
@@ 2.3 Feature Selection Methods
Feature selection is another simple and intuitive method for removing redundant features. Feature selection methods can be applied to bag-of-words representations, where each unique word is treated as a distinct feature, for emotion recognition [5] and sentiment classification [25] tasks to reduce the number of terms and save storage and memory space. Also, feature selection methods can be applied for text classification to select the most representative word [20]. However, the role of feature selection for word embedding dimension reduction is under-explored. This paper shows that a straightforward and intuitive feature selection method can achieve effective dimension reduction. To our knowledge, applying the feature selection method for word embedding dimension reduction is novel.
@@ 3.2 Post-processing
To improve the quality of word embedding in downstream applications, performing post-processing techniques on pre-trained word embeddings can be crucial. The post-processing algorithm (PPA) [15] is an effective post-processing method to improve the isotropy of word representations by removing the top principal components of all words. We explain the details of PPA in Algorithm  1 [15].
The existing effective PCA-based method [19] applies the PPA before and after dimension reduction to enhance both the original and the dimension-reduced word vectors. In our method, we only apply PPA before dimension reduction, which makes our method simpler. Applying PCA-based post-processing to the selected subset after feature selection may disrupt the previous results because PCA and the feature selection criteria can have different objectives. Specifically, PCA transforms features to maximize variance, while feature selection methods usually select features based on their correlation with labels. Also, the selected feature subset may lose crucial information for PCA to find valuable principal components. As an enhancement of original word embeddings before dimension reduction, we also make PPA an optional choice. The reason is that anisotropy may not always be harmful [1] and using PPA depends on the pre-trained word embeddings and specific application scenarios. Specifically, word embeddings trained on less data may be vulnerable to noise. Applying PPA helps mitigate some of the noise and offers better performance. On the other hand, using PPA may result in the loss of crucial and intricate information related to word representation, which can harm performance. Word embeddings trained on relatively small data benefit from uniformly distributed word embeddings, while complex contextual tasks may require anisotropy to capture detailed information.
@@ 3.3 Feature Extraction
$\text{Sim}(a,b)$ $=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$ (1)
@@
where $\mathbf{a}$ and $\mathbf{b}$ are embeddings of the two words. The cosine similarity score can be viewed as the sum of the normalized results of element-wise products. Each pair of elements (i.e., each dimension) has a linear impact on the cosine similarity score of that pair of word vectors. Therefore, the normalized results of element-wise products can be viewed as features to predict the similarity score. The feature we use for each dimension is:
@@
where $f_{i}$ is the feature of the $i^{th}$ dimension extracted from the two word embeddings.
@@ 3.4 Feature Selection
where $b=1,2,...B-1$. The threshold $t$ partitions the $i^{th}$ feature space into the left subset $S_{L,t}^{i}$ and the right subset $S_{R,t}^{i}$. The MSEs for regression $R_{L,t}^{i}$ and $R_{R,t}^{i}$ are calculated separately in the two subsets. Then, the RFT loss is defined as the weighted sum of the two MSEs:
@@
where $N_{L,t}^{i}$ and $N_{R,t}^{i}$ denote the number of samples in each subset and $N$ is the total number of samples. The RFT loss of each feature dimension is defined as the minimal RFT loss among all the candidate thresholds:
@@
$r_{i}=\rho_{R(f_{i}),R(y)}=\frac{\text{cov}(R(f_{i}),R(y))}{\sigma_{R(f_{i})}\sigma_{R(y)}}.$ (6)
The higher the correlation coefficient, the better the feature dimension is. Finally, the Spearman’s rank correlation coefficients of all the feature dimensions are then ranked in descending order, and the top $K$ dimensions are selected.
@@ 4 Experiments
We evaluate our method by applying it to several pre-trained word embeddings. The word similarity datasets provide weak supervision to guide our feature selection module. The dimension-reduced word embeddings are used for word similarity tasks and downstream tasks. Then, we compare the results of our method with the original pre-trained word embeddings, the dimension-reduced word embeddings from the PCA-based method called Algo [19], and the results from the deep learning-based method called EmbedTextNet [8].
@@ 4.2 Word Similarity Datasets
We first evaluate our methods on twelve word similarity datasets. Since we choose word similarity as the guidance for feature selection, we apply 5-fold cross-validation to each word similarity dataset to train and test our method. The reported results are the average Spearman’s rank correlation coefficients from the five folds. We evaluate the pre-trained word embeddings and all the methods on the same cross-validation fold and take the average for a fair comparison. All the results we report average over five different cross-validation trials in all the experiments. We set the number of bins in RFT to 4, the default and recommended value from the RFT model, and the threshold in the PPA to 7, the same as in the PCA-based model.
@@
WordFS-P-woP-150D(ours) 66.79 74.94 60.68 63.44 69.76 36.16 37.93 30.30 62.49 53.69 66.78 52.27 56.27
WordFS-P-wP-150D(ours) 72.94 75.38 64.38 64.24 70.46 45.01 42.47 34.88 68.53 60.58 70.25 50.67 59.98
@@
Table 1 shows the results across twelve word similarity datasets of different dimension reduction methods reduced from 300-dimensional Glove word embeddings to 150, 100, and 50 dimensions, respectively. In the table, WordFS-P and WordFS-S represent our proposed methods based on RFT and similarity feature selection methods, respectively. P represents PPA, and wP and woP represent PPA and without PPA. Algo is the method proposed in [19], which utilizes PPA before and after PCA. The last column is the average of the previous columns. For the dimension-reduced vectors with dimensions of 50, 100, and 150, our methods outperform the original word embeddings and other methods in most datasets across these reduced dimensions. Our WordFS-S-wP method achieves the highest average score when the dimension is reduced to 50 and 150 and is close to the highest, which is our WordFS-P-wP method, with a difference of 0.03 when reduced to 100. Both feature selection-based approaches perform better than the existing methods. As expected, the WordFS-S-wP method performs better than the WordFS-P-wP method in this task. Our WordFS-S-wP method outperforms the best performance from existing methods with an average improvement of 0.61, 1.97, and 3.69 in terms of Spearman’s rank correlation coefficients when reducing to 150, 100, and 50 dimensions, respectively. Our method can perform better than the original 300-dimensional word embeddings, even when reduced to 100 dimensions.
@@
Table 2 shows the results of various methods for dimension reduction on the aggregated dataset. The annotations in the tables are the same as in the previous one. Our WordFS-S methods perform best among all the experiments with different pre-trained word embeddings in other dimensions. And there is a large gap between our methods and the existing methods. Our approaches significantly benefit this task compared to the PCA-based method, particularly in lower dimensions where the gap widens. The performance of the EmbedTextNet model improves at lower dimensions but still lags behind our WordFS-S methods. It is worth noting that even when reduced to 50 dimensions, our WordFS-S-wP method achieves better performance compared to the original 300-dimensional word embeddings. Our experiment with three different pre-trained word embeddings shows that our method achieved the best performance for the Glove and Fasttext embeddings when post-processing is applied. However, for the Word2vec embeddings, better results are obtained without post-processing. This is reasonable because the power of the post-processing method may depend on the pre-trained word embeddings and the specific application scenario. The Word2vec embeddings we used were pre-trained on the most extensive data among the three. It implies that the embeddings may be more resilient to noise and capable of capturing more nuanced information. However, applying PPA can potentially result in the loss of information and cause a drop in performance.

### v2

@@ Abstract
As a fundamental task in natural language processing, word embedding converts each word into a representation in a vector space. A challenge with word embedding is that as the vocabulary grows, the vector space’s dimension increases, which can lead to a vast model size. Storing and processing word vectors are resource-demanding, especially for mobile edge-devices applications. This paper explores word embedding dimension reduction. To balance computational costs and performance, we propose an efficient and effective weakly-supervised feature selection method named WordFS. It has two variants, each utilizing novel criteria for feature selection. Experiments on various tasks (e.g., word and sentence similarity and binary and multi-class classification) indicate that the proposed WordFS model outperforms other dimension reduction methods at lower computational costs. We have released the code for reproducibility along with the paper  [^1] https://github.com/jintangxue/WordFS.
@@ 1 Introduction
In recent years, large language models (LLMs) have made a breakthrough in natural language processing (NLP). These models have revolutionized the understanding and generation of human language tasks such as question-answering, machine translation, and text summarization [29, 30, 17], etc. As a fundamental component in language models, word embedding represents words as vectors in a continuous, high-dimensional space [28]. Word vectors capture semantic and syntactic meanings of words, providing word relationships for various downstream tasks. Static word embedding methods, such as Glove [18], Word2Vec [14], and Fasttext [2], assign a fixed vector to each word. They convert input words to corresponding vectors for further processing and model training. Contextual word embedding methods, such as ELMo [19] and BERT [4], leverage deep-learning models to generate word vectors based on the context. A word can have a different contextual word embedding in a different sentence, allowing the model to capture subtle differences in the meaning of the same word in various contexts.
@@
Traditional unsupervised PCA-based methods are known for their efficiency and interpretability. In contrast, supervised deep-learning-based methods are inefficient and lack interpretability. For instance, autoencoders take longer training and inference time in sentence embedding dimension reduction than traditional methods [32]. Semi-supervised feature selection methods can also be used for dimension reduction. Since word vectors can be viewed as extracted features for each word, feature selection methods have the potential to provide a straightforward yet effective way to reduce the word dimension. It could help balance unsupervised PCA-based methods and resource-intensive deep-learning-based methods.
This paper investigates using a small subset of labeled word similarity pairs to develop a weakly supervised dimension reduction method called WordFS (i.e., word dimension reduction with Feature Selection). We demonstrate that one can achieve dimension reduction effectively by supervising a limited number of word similarity pairs. Note that word similarity is typically used in evaluation tasks but is rare in word embedding dimension reduction, a key novelty of this work. The proposed WordFS method consists of three stages: 1) post-processing, 2) feature extraction, and 3) weakly-supervised feature selection. We apply WordFS to other downstream tasks to show its generalizability. Experimental results show that WordFS outperforms existing methods in word similarity and various tasks while achieving much lower computational costs.
@@
  • We show the effectiveness and efficiency of our approach on various downstream tasks, including sentence similarity and classification tasks. Our method generally outperforms the existing techniques while being more straightforward and efficient.
@@ 2.1 Matrix Decomposition Techniques
Post-processing methods help enhance word embeddings in dimension reduction. Effective construction of lower dimensional word embeddings can be achieved by combining post-processing algorithms (PPAs) based on reducing the anisotropy [16] with the PCA method [20]. The dimension-reduced vectors can perform similarly or even better than the original pre-trained word embeddings, outperforming most unsupervised methods. [9] extends this method to contextual embeddings by adding in additional geodesic distance information via the Isomap algorithm [25]. PCA also performs well as an unsupervised dimension reduction method for pre-trained sentence embeddings [32].
2.2 Non-Linear Dimensionality Reduction Techniques
Non-linear dimensionality reduction techniques like uniform manifold approximation and projection (UMAP) and t-distributed stochastic neighbor embedding (t-SNE) have been used on dimension reduction tasks. UMAP is a dimensionality reduction method based on manifold theory and topological data analysis. It constructs a topological representation of high-dimensional data by approximating local manifolds and stitching together their fuzzy simplicial sets. However, UMAP’s performance in reducing the dimensionality of word embeddings is not satisfactory. t-SNE is a statistical method based on stochastic neighbor embedding. It assigns each data point a location in a two- or three-dimensional map for visualizing high-dimensional data. Applying t-SNE to a large corpus while keeping many dimensions requires a significant amount of time and memory, making it unsuitable for this task.
@@ 2.4 Feature Selection Methods
Feature selection is another simple and intuitive method for removing redundant features. Feature selection methods can be applied to bag-of-words representations, where each unique word is treated as a distinct feature, for emotion recognition [5] and sentiment classification [26] tasks to reduce the number of terms and save storage and memory space. Also, feature selection methods can be applied for text classification to select the most representative word [21]. However, the role of feature selection for word embedding dimension reduction is under-explored. This paper shows that a straightforward and intuitive feature selection method can effectively reduce dimension. We believe applying the feature selection method for word embedding dimension reduction is novel.
@@ 3.2 Post-processing
To improve the quality of word embedding in downstream applications, performing post-processing techniques on pre-trained word embeddings can be crucial. The post-processing algorithm (PPA) [16] is an effective post-processing method to improve the isotropy of word representations by removing the common mean vector and the top principal components of all words. Details of PPA are shown in Algorithm  1 [16]. First, the mean vector is computed from all embeddings in the vocabulary and then subtracted from each embedding. Next, the PCA components are derived from the adjusted mean matrix, with the top-D components being chosen. Finally, the projections on the top-D principal components are removed from each word embedding.
The existing effective PCA-based method [20] applies the PPA before and after dimension reduction to enhance both the original and the dimension-reduced word vectors. In our method, we only apply PPA before dimension reduction, which makes our method simpler. Applying PCA-based post-processing to the selected subset after feature selection may disrupt the previous results because PCA and the feature selection criteria can have different objectives. Specifically, PCA transforms features to maximize variance, while feature selection methods usually select features based on their correlation with labels. Also, the selected feature subset may lose crucial information for PCA to find valuable principal components. As an enhancement of original word embeddings before dimension reduction, we also make PPA an optional choice. The reason is that anisotropy may not always be harmful [1], and using PPA depends on the pre-trained word embeddings and specific application scenarios. Specifically, word embeddings trained on less data may be vulnerable to noise. Applying PPA helps mitigate some of the noise and offers better performance. On the other hand, using PPA may result in the loss of crucial and intricate information related to word representation, which can harm performance. Word embeddings trained on relatively small data benefit from uniformly distributed word embeddings, while complex contextual tasks may require anisotropy to capture detailed information.
@@ 3.3 Feature Extraction
$\text{Sim}(\mathbf{a},\mathbf{b})$ $=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$ (1)
@@
where $\mathbf{a}$ and $\mathbf{b}$ are embeddings of the two words, and n represents the number of elements in a word embedding. $a_{i}$ and $b_{i}$ are the $i$th elements of $\mathbf{a}$ and $\mathbf{b}$, respectively. The cosine similarity score can be viewed as the sum of the normalized results of element-wise products. Each pair of elements (i.e., each dimension) has a linear impact on the cosine similarity score of that pair of word vectors. Therefore, the normalized results of element-wise products can be viewed as features to predict the similarity score. The feature we use for each dimension is:
@@
where $f_{i}$ is the feature of the $i^{th}$ dimension extracted from thecorresponding elements $a_{i}$ and $b_{i}$ of the two word embeddings.
@@ 3.4 Feature Selection
where $b=1,2,...B-1$. A threshold $t$ partitions the $i^{th}$ feature space into the left subset $S_{L,t}^{i}$ and the right subset $S_{R,t}^{i}$. The MSEs for regression $R_{L,t}^{i}$ and $R_{R,t}^{i}$ are separately calculated in the two subsets by using the mean of target values as the estimated regression value of all samples. Then, the RFT loss is defined as the weighted sum of the two MSEs:
@@
where $N_{L,t}^{i}$ and $N_{R,t}^{i}$ denote the number of samples in each subset and $N$ is the total number of samples. The RFT loss of the $i$th feature dimension is defined as the minimal RFT loss among all the candidate thresholds:
@@
$r_{i}=\rho_{R(f_{i}),R(y)}=\frac{\text{cov}(R(f_{i}),R(y))}{\sigma_{R(f_{i})}\sigma_{R(y)}},$ (6)
where $R(\cdot)$ represents the rank of given raw scores. The higher the correlation coefficient, the better the feature dimension is. Finally, the Spearman’s rank correlation coefficients of all the feature dimensions are then ranked in descending order, and the top $K$ dimensions are selected.
@@ 4 Experiments
We evaluate our method by applying it to several pre-trained word embeddings. The word similarity datasets provide weak supervision to guide our feature selection module. The dimension-reduced word embeddings are used for word similarity tasks and downstream tasks. Then, we compare the results of our method with the original pre-trained word embeddings, the reduced word embeddings from UMAP [13], the dimension-reduced word embeddings from the PCA-based method called Algo [20], and the results from the deep learning-based method called EmbedTextNet [8]. The performance of different datasets can vary due to their specific requirements and emphasis. Finding a general dimensionality reduction method that performs well across all datasets without fine-tuning is challenging. Therefore, we focus on comparing the average performance for each task. The method that achieves better performance on average is considered better.
@@ 4.2 Word Similarity Datasets
We first evaluate our methods on twelve word similarity datasets based on all the three word embeddings. In later experiments, we will assess the generalizability of our WordFS model across various embeddings on multiple downstream tasks. Since we choose word similarity as the guidance for feature selection, we apply 5-fold cross-validation to each word similarity dataset to train and test our method. The reported results are the average Spearman’s rank correlation coefficients from the five folds. We evaluate the pre-trained word embeddings and all the methods on the same cross-validation fold and take the average for a fair comparison. All the results we report average over five different cross-validation trials in all the experiments. We set the number of bins in RFT to 4, the default and recommended value from the RFT model, and the threshold in the PPA to 7, the same as in the PCA-based model.
@@
WordFS-P-woP-150D(ours) 66.79 74.94 60.68 63.44 69.83 36.16 37.93 30.26 62.49 53.69 66.78 52.27 56.27
WordFS-P-wP-150D(ours) 73.63 75.38 64.38 64.24 70.37 45.01 42.47 34.88 68.53 60.58 70.25 50.54 60.02
@@
Word2vec-300D 73.42 77.01 68.52 66.73 71.65 53.39 44.19 49.69 68.92 61.42 76.87 53.80 63.80
UMAP-150D 36.38 50.43 42.53 26.17 35.13 26.00 22.32 6.25 42.09 32.59 51.84 17.37 32.42
Algo-150D 78.91 78.23 63.46 65.86 74.93 51.43 42.53 41.13 68.25 60.28 75.11 48.34 62.37
EmbedTextNet-150D 72.28 76.33 63.09 65.28 72.33 52.27 43.53 44.47 66.44 57.95 75.04 48.89 61.49
WordFS-P-woP-150D(ours) 68.16 76.16 63.35 63.39 70.79 51.60 45.32 55.99 67.68 58.04 72.44 48.14 61.75
WordFS-P-wP-150D(ours) 66.80 76.77 63.43 63.45 69.99 52.00 46.57 50.77 66.81 56.57 72.64 48.75 61.21
WordFS-S-woP-150D(ours) 75.71 75.93 64.22 63.35 71.63 51.86 46.95 51.89 69.58 61.30 72.85 53.07 63.19
WordFS-S-wP-150D(ours) 74.80 76.73 61.72 63.74 70.55 52.59 46.77 49.56 68.79 58.41 74.73 51.99 62.53
UMAP-100D 33.18 51.38 42.27 27.97 33.24 26.86 23.03 8.13 42.56 31.03 53.53 11.16 32.03
Algo-100D 80.06 76.08 63.86 63.56 72.26 48.21 39.76 30.78 67.49 59.81 73.40 39.65 59.58
EmbedTextNet-100D 74.11 75.73 62.21 63.14 71.10 50.33 41.89 40.12 66.10 59.38 72.89 45.34 60.19
WordFS-P-woP-100D(ours) 62.45 74.14 57.35 61.32 69.55 50.01 45.18 52.28 64.82 55.96 71.10 44.29 59.04
WordFS-S-woP-100D(ours) 65.42 74.45 60.81 60.28 66.60 49.75 45.70 52.60 68.49 57.76 70.23 50.48 60.21
UMAP-50D 28.38 50.06 42.88 27.44 35.31 25.90 22.21 9.78 44.41 34.64 54.23 16.77 32.67
Algo-50D 73.87 70.36 62.29 54.36 66.97 41.89 32.36 22.72 59.75 53.79 66.67 37.28 53.53
EmbedTextNet-50D 72.49 73.10 66.33 59.24 62.16 46.30 36.71 26.65 63.99 55.74 72.62 35.26 55.88
WordFS-P-woP-50D(ours) 49.65 69.72 53.24 55.01 63.96 44.47 43.12 41.79 57.92 47.89 65.01 34.51 52.19
WordFS-S-woP-50D(ours) 54.68 70.87 51.66 54.62 55.86 44.65 43.52 47.26 61.96 49.05 65.16 38.93 53.18
Fasttext-300D 77.08 79.01 69.84 70.57 82.68 52.23 45.11 43.86 70.63 64.65 79.66 48.92 65.35
UMAP-150D 56.94 59.32 46.98 43.50 49.67 33.78 29.16 32.07 40.20 29.54 54.10 20.65 41.33
Algo-150D 92.63 80.59 71.32 70.74 87.97 51.11 45.04 45.03 73.18 68.03 77.21 49.31 67.68
EmbedTextNet-150D 77.54 78.50 68.39 69.78 82.10 52.03 45.71 38.11 68.01 62.98 74.77 51.57 64.12
WordFS-P-woP-150D(ours) 70.45 79.52 65.50 67.05 80.15 51.29 46.59 39.11 69.56 60.91 77.27 51.88 63.27
WordFS-P-wP-150D(ours) 76.17 80.13 67.26 70.29 79.83 54.58 47.44 38.84 72.93 66.76 78.53 55.68 65.70
WordFS-S-woP-150D(ours) 69.54 80.39 64.20 69.73 75.89 53.47 47.43 44.62 74.32 66.94 77.07 52.69 64.69
WordFS-S-wP-150D(ours) 73.42 80.89 66.50 68.87 78.78 53.94 49.14 46.14 74.99 71.09 77.40 53.71 66.24
UMAP-100D 55.58 59.31 46.16 43.69 41.73 33.21 28.80 32.09 39.17 29.49 53.00 19.40 41.33
Algo-100D 89.65 78.80 71.86 66.88 87.66 49.19 42.30 39.96 70.33 64.11 74.89 45.88 65.13
EmbedTextNet-100D 81.65 77.74 68.26 66.93 81.72 51.34 40.62 41.30 69.21 62.20 75.68 46.99 63.64
WordFS-P-wP-100D(ours) 70.91 77.40 63.64 66.07 76.81 52.08 46.48 35.23 71.05 65.77 74.08 52.72 62.69
WordFS-S-wP-100D(ours) 75.71 78.33 65.38 66.44 76.89 51.25 47.24 43.04 72.63 67.27 75.96 56.67 64.73
UMAP-50D 55.12 58.52 46.70 43.67 40.37 33.63 29.38 31.36 40.87 31.62 54.07 21.06 40.53
Algo-50D 79.37 72.74 70.66 61.06 77.50 45.47 35.79 37.48 59.66 52.72 67.79 43.79 58.67
EmbedTextNet-50D 74.80 72.39 69.21 60.69 73.40 45.50 36.04 42.57 70.13 62.00 76.17 34.52 59.78
WordFS-P-wP-50D(ours) 61.31 70.80 54.88 58.09 67.07 46.60 44.66 26.71 64.36 57.75 66.41 50.27 55.74
WordFS-S-wP-50D(ours) 67.94 71.97 62.42 60.33 67.92 45.53 45.48 41.89 67.41 59.24 70.99 54.36 59.62
@@
Table 1 shows the results across twelve word similarity datasets of different dimension reduction methods reduced from 300-dimensional Glove word embeddings to 150, 100, and 50 dimensions, respectively. In the table, WordFS-P and WordFS-S represent our proposed methods based on RFT and similarity feature selection methods, respectively. P represents PPA, and wP and woP represent PPA and without PPA. Algo is the method proposed in [20], which utilizes PPA before and after PCA. The last column is the average of the previous columns. The UMAP results are significantly lower compared to other methods. This is understandable because UMAP is a general dimension reduction method and is not specifically designed to reduce the dimension of word embeddings, unlike other methods. For the dimension-reduced vectors with dimensions of 50, 100, and 150 based on the Glove embedding, our methods outperform the original Glove word embeddings and other methods in most datasets across these reduced dimensions. Our WordFS-S-wP method achieves the highest average score when the dimension is reduced to 50 and 150 and is close to the highest, which is our WordFS-P-wP method, with a difference of 0.03 when reduced to 100. Both feature selection-based approaches perform better than the existing methods. As expected, the WordFS-S-wP method performs better than the WordFS-P-wP method in this task. Our WordFS-S-wP method outperforms the best performance from existing methods with an average improvement of 0.61, 1.97, and 3.69 in terms of Spearman’s rank correlation coefficients when reducing to 150, 100, and 50 dimensions, respectively. Our method can perform better than the original 300-dimensional word embeddings, even when reduced to 100 dimensions. For the dimension-reduced vectors based on the Word2vec embeddings, applying PPA may decrease performance, as the word embeddings are pre-trained on larger data, containing subtle linguistic connections. Our method can perform highest when the word vector dimension is reduced to 150 and 100 and slightly lower when reduced to 50. For the Fasttext word embeddings, although our method may not be the best, it consistently performs very close to the highest performance.
@@
Table 2 shows the results of various methods for dimension reduction on the aggregated dataset. The annotations in the tables are the same as in the previous one. The performance of UMAP is still very low. Our WordFS-S methods perform best among all the experiments with different pre-trained word embeddings in other dimensions. And there is a large gap between our methods and the existing methods. Our approaches significantly benefit this task compared to the PCA-based method, particularly in lower dimensions where the gap widens. The performance of the EmbedTextNet model improves at lower dimensions but still lags behind our WordFS-S methods. It is worth noting that even when reduced to 50 dimensions, our WordFS-S-wP method achieves better performance compared to the original 300-dimensional word embeddings. Our experiment with three different pre-trained word embeddings shows that our method achieved the best performance for the Glove and Fasttext embeddings when post-processing is applied. However, for the Word2vec embeddings, better results are obtained without post-processing. This is reasonable because the power of the post-processing method may depend on the pre-trained word embeddings and the specific application scenario. The Word2vec embeddings we used were pre-trained on the most extensive data among the three. It implies that the embeddings may be more resilient to noise and capable of capturing more nuanced information. However, applying PPA can potentially result in the loss of information and cause a drop in performance.